# Symplectic mapping class relations from pencil pairs

### v1

@@ 4.1. Projective models
Let $(Z,\operatorname{\mathcal{L}})$ be a polarized variety having $\dim_{\mathbb{C}}Z=n$, let $\mathcal{M}\subset\mathbb{P}^{N}$ be a projective variety of $\dim_{\mathbb{C}}\mathcal{M}=m\geq n$, and let deg=($\deg=(d_{1},\dots,d_{m-n})\in\mathbb{Z}^{m-n}_{>0}$. We say that $(\mathcal{M},\deg)$ is a projective model for $(Z,\operatorname{\mathcal{L}})$ if there is a a holomorphic section $\boldsymbol{\mathfrak{s}}_{\mathcal{M}}$ of $\mathcal{O}_{\deg}=\oplus_{1}^{m-n}\mathcal{O}_{d_{j}}\rightarrow\mathbb{P}^{N}$ such that
  1. (1)
    there is a biholomorphism $\phi:Z\rightarrow Z_{\mathcal{M}}=\{\boldsymbol{\mathfrak{s}}_{\mathcal{M}}=0\}\cap\mathcal{M}$, and
  2. (2)
    $\phi^{\ast}\mathcal{O}_{\deg_{1}}=\operatorname{\mathcal{L}}\rightarrow Z$.
Lemma 4.1.2.
Let $(X^{\pm},\operatorname{\mathcal{L}}^{\pm})$ be a pair of ample polarized varieties. Suppose that there are holomorphic sections $\boldsymbol{\mathfrak{s}}^{\pm}_{1}$ of $\operatorname{\mathcal{L}}^{\pm}$ for which
@@
    the $(Z^{\pm},\operatorname{\mathcal{L}}^{\pm})$ have a common projective model $(\mathcal{M},\deg)$.
Then there are pencils for the ±$\operatorname{\mathcal{L}}^{\pm}\rightarrow X^{\pm}$ and a $\phi$ for which $(X^{\pm},\operatorname{\mathcal{L}}^{\pm},\phi)$ is a pencil pair.
@@
By the definition of projective models there are sections $\boldsymbol{\mathfrak{s}}^{\pm}_{\mathcal{M},1}$ of $\mathcal{O}_{\deg}$ with zero loci $Z_{\mathcal{M}}^{\pm}\subset\mathcal{M}$ and biholomorphisms $\phi^{\pm}:Z^{\pm}_{1}\rightarrow Z_{\mathcal{M}}^{\pm}$ for which ±$\operatorname{\mathcal{L}}^{\pm}|_{Z_{1}^{\pm}}=\phi^{\ast}\mathcal{O}_{\deg_{1}}$. Choose sections $\boldsymbol{\mathfrak{s}}^{\pm}_{0}$ of the $\operatorname{\mathcal{L}}^{\pm}$ so that the $(\boldsymbol{\mathfrak{s}}^{\pm}_{0},\boldsymbol{\mathfrak{s}}^{\pm}_{1})$ are pencils for the $(X^{\pm},\operatorname{\mathcal{L}}^{\pm})$ with base loci $B^{\pm}=\{\boldsymbol{\mathfrak{s}}^{\pm}_{0}=\boldsymbol{\mathfrak{s}}^{\pm}_{1}=0\}\subset Z^{\pm}_{1}$. By the existence of the projective models, there are sections $\boldsymbol{\mathfrak{s}}_{\mathcal{M},0}^{\pm}$ of $\mathcal{O}_{\deg_{1}}\rightarrow\mathbb{P}^{N}$ such that $\phi^{\pm}B^{\pm}\subset\mathcal{M}$ is cut out by the section $\boldsymbol{\mathfrak{s}}^{\pm}_{\mathcal{M},1}\oplus\boldsymbol{\mathfrak{s}}_{\mathcal{M},0}^{\pm}$ of $\mathcal{O}_{\deg}\oplus\mathcal{O}_{1}\rightarrow\mathbb{P}^{N}$.
Let $\boldsymbol{\mathfrak{s}}^{T}_{\mathcal{M},1}\oplus\boldsymbol{\mathfrak{s}}^{T}_{\mathcal{M},0}$ be a $T\in[-1,1]$ parameter family of sections of $\mathcal{O}_{\deg}\oplus\mathcal{O}_{1}\rightarrow\mathbb{P}^{m}$ for which $\boldsymbol{\mathfrak{s}}^{\pm 1}_{\mathcal{M}}=\boldsymbol{\mathfrak{s}}_{\mathcal{M}}^{\pm}$, $\boldsymbol{\mathfrak{s}}^{\pm 1}_{\mathcal{M},0}=\boldsymbol{\mathfrak{s}}_{\mathcal{M},0}^{\pm}$ and such that the zero loci of the $\boldsymbol{\mathfrak{s}}^{T}_{\mathcal{M},1}$ and $\boldsymbol{\mathfrak{s}}^{T}_{\mathcal{M}}\oplus\boldsymbol{\mathfrak{s}}^{T}_{\mathcal{M},0}$ are transversely cut out for all $T$. The transversality conditions can be satisfied using the fact that failure of transversality is a $\operatorname{codim}_{\mathbb{C}}\geq 1$ condition in the projective spaces of sections $\mathbb{P}(H^{0}(\mathcal{O}_{\deg}\oplus\mathcal{O}_{1}))$. A parametric Darboux-Moser-Weinstein argument gives us symplectomorphism
$\psi:(\phi^{+}Z^{\pm}_{1},\omega_{\mathcal{O}_{1}},\phi^{+}B^{+}=\{\boldsymbol{\mathfrak{s}}^{+}_{\mathcal{M}}\oplus\boldsymbol{\mathfrak{s}}^{+}_{\mathcal{M},0}=0\})\rightarrow(\phi^{-}Z^{\pm}_{1},\omega_{\mathcal{O}_{1}},\phi^{+}B^{-}=\{\boldsymbol{\mathfrak{s}}^{-}_{\mathcal{M}}\oplus\boldsymbol{\mathfrak{s}}^{-}_{\mathcal{M},0}=0\}).$
So $\phi=(\phi^{-})^{-1}\psi\phi^{+}$ a symplectomorphism $(Z^{+}_{1},\omega_{\operatorname{\mathcal{L}}^{+}},B^{+})\rightarrow(Z^{-}_{1},\omega_{\operatorname{\mathcal{L}}^{-}},B^{-})$. ∎
Definition 4.1.3.
In the above notation, $(X^{\pm},\operatorname{\mathcal{L}}^{\pm},\phi)$ is determined by the projective model $(\mathcal{M},\deg)$.
@@ 4.2. Cabling pencil pairs
The following theorem allows us to determine an infinite family of pencil pairs from a single pencil pair when it is determined by a projective model.
Lemma 4.2.1.
Let $(X^{\pm},\operatorname{\mathcal{L}}^{\pm},\phi^{\pm})$ be determined by the projective model $(\mathcal{M},\deg=(d_{1},\dots,d_{m-n}))$. Then for all $k\in\mathbb{Z}_{\geq 1}$ there are $\phi_{k}$ for which there exists pencil pairs $(X^{\pm},(\operatorname{\mathcal{L}}^{\pm})^{\otimes k},\phi_{k})$ determined by the projective model $(\mathcal{M},\deg_{k}=(kd_{1},d_{2},\dots,d_{m-n}))$.
@@
To start we study the local topology of divisors associated to powers of a line bundle. Let $Z\subset X$ be the zero locus of some transversely cut out section $\boldsymbol{\mathfrak{s}}$ of an ample $\operatorname{\mathcal{L}}\rightarrow X$. Consider sections $\boldsymbol{\mathfrak{s}}_{k}$ of $\operatorname{\mathcal{L}}^{\otimes k}$ the form $\boldsymbol{\mathfrak{s}}_{k}=\boldsymbol{\mathfrak{s}}^{\otimes k}+\mathfrak{t}$ with $\mathfrak{t}$ a $\mathcal{C}^{1}$ small section of $\operatorname{\mathcal{L}}^{\otimes k}$. By the smallness of $\mathfrak{t}$, the zero locus $Z_{k}$ of $\boldsymbol{\mathfrak{s}}_{k}$ is contained in an arbitrarily small tubular neighborhood of $Z$ in $X$. Writing $B_{k}=Z_{k}\cap Z=\{\mathfrak{t}=0\}\subset Z$, we see that $Z\setminus B_{k}$ is an affine variety since it is the complement of a hyperplane section in a projective variety. Hence $Z\setminus B_{k}$ is Stein. So by [Si76, Corollary 2] it has a neighborhood in $X$ whose projection onto $Z\setminus B_{k}$ is holomorphic. Possibly after multiplying $\mathfrak{t}$ by a small constant (which has no effect on $B_{k}$), we can assume that $Z_{k}\setminus B_{k}$ is contained in this neighborhood. The projection of this tubular neighborhood to $Z\setminus B_{k}$ restricted to $Z_{k}\setminus B_{k}$ gives a $k$-fold covering $\pi_{k}:Z_{k}\setminus B_{k}\rightarrow Z\setminus B_{k}$. By compactifying, we obtain a $k$-fold branched covering $\overline{\pi}_{k}:Z_{k}\rightarrow Z$ with branch locus $B_{k}$.
We claim that the isomorphism type of $(Z_{k},\operatorname{\mathcal{L}}^{\otimes k})$ depends only on $(Z,B_{k})$. The section $\mathfrak{t}$ trivializes $\operatorname{\mathcal{L}}^{\otimes k}$ over $Z_{k}\setminus B_{k}$ through which a holonomy morphism $\psi:\pi_{1}(Z\setminus Z_{k})\rightarrow\mathbb{S}^{1}$ is determined. Covering space theory tells us that the cover $\pi_{k}$ is smoothly classified by $\psi$ and so $\overline{\pi}_{k}$ is classified by $\psi$.
Since $\pi_{k}$ is holomorphic and a covering map, it is a local biholomorphism. Identifying $Z_{k}$ with the closure of $Z_{k}\setminus B_{k}$ in $X$, we see that the biholomorphism type of $Z_{k}$ is completely determined by the pair $(Z,B_{k})$. Observing that $B_{k}\subset Z_{k}$ is the zero locus $\boldsymbol{\mathfrak{s}}|_{Z_{k}}:Z_{k}\rightarrow\operatorname{\mathcal{L}}$ and applying the line-bundle/divisor correspondence, we see that the isomorphism type of $(Z_{k},\operatorname{\mathcal{L}})$ depends only on $(Z,B_{k})$. Hence the desired claim follows for $(Z_{k},\operatorname{\mathcal{L}}^{\otimes k})$.
@@
It follows that is $(\mathcal{M},\deg)$ is a projective model for $(Z,\operatorname{\mathcal{L}})$, then $(\mathcal{M},\deg_{k})$ is a projective model for $(Z_{k},\operatorname{\mathcal{L}}^{\otimes k})$. Applying $\pm$ signs, we see that if $(X^{\pm},\operatorname{\mathcal{L}}^{\pm},\phi)$ are determined by a projective model, then we can use Lemma 4.1.2 to find $\phi_{k}$ for which the $(X^{\pm},(\operatorname{\mathcal{L}}^{\pm})^{\otimes k},\phi_{k})$ are pencil pairs determined by the projective models $(\mathcal{M},\deg_{k})$. ∎

### v2

@@ 4.1. Projective models
Let $(Z,\operatorname{\mathcal{L}})$ be a polarized variety of $\dim_{\mathbb{C}}Z=n$, $\mathcal{M}\subset\mathbb{CP}^{N}$ be a projective variety of $\dim_{\mathbb{C}}\mathcal{M}=m>n$, $\vec{d}=(d_{1},\dots,d_{m-n})\in\mathbb{Z}^{m-n}_{>0}$, and write $\mathcal{O}_{\vec{d}}=\oplus_{j=1}^{m-n}\mathcal{O}_{d_{j}}$. We say that $(\mathcal{M},\vec{d})$ is a projective model for $(Z,\operatorname{\mathcal{L}})$ if there is a holomorphic section of $\boldsymbol{\mathfrak{s}}_{\mathcal{M}}$ of $\mathcal{O}_{\vec{d}}\to\mathbb{CP}^{N}$ and a biholomorphism
$\phi:Z\rightarrow Z_{\mathcal{M}},\quad Z_{\mathcal{M}}=\{\boldsymbol{\mathfrak{s}}_{\mathcal{M}}=0\}\cap\mathcal{M}$
for which $\phi^{\ast}\mathcal{O}_{d_{1}}=\operatorname{\mathcal{L}}$.
Definition 4.1.2.
Let $(X^{\pm},\operatorname{\mathcal{L}}^{\pm})$ be a pair of very ample polarized varieties. Suppose we have holomorphic sections $\boldsymbol{\mathfrak{s}}^{\pm}_{1}\in H^{0}(X^{\pm},\operatorname{\mathcal{L}}^{\pm})$ for which
@@
    the polarized varieties $(Z_{1}^{\pm},\operatorname{\mathcal{L}}^{\pm})$ have a common projective model $(\mathcal{M},\vec{d})$.
Then we say that the $(X^{\pm},\operatorname{\mathcal{L}}^{\pm})$ have divisors with a common projective model.
Lemma 4.1.3.
Suppose that the $(X^{\pm},\operatorname{\mathcal{L}}^{\pm})$ have divisors with a common projective model. Then there are pencils for the ±$\operatorname{\mathcal{L}}^{\pm}\rightarrow X^{\pm}$ and a $\phi$ for which $(X^{\pm},\operatorname{\mathcal{L}}^{\pm},\phi)$ is a pencil pair.
@@
By the definition of projective models there are sections $\boldsymbol{\mathfrak{s}}^{\pm}_{\mathcal{M},1}$ of $\mathcal{O}_{\vec{d}}$ with zero loci $Z_{\mathcal{M}}^{\pm}\subset\mathcal{M}$ and biholomorphisms $\phi^{\pm}:Z^{\pm}_{1}\rightarrow Z_{\mathcal{M}}^{\pm}$ for which ±$\operatorname{\mathcal{L}}^{\pm}|_{Z_{1}^{\pm}}=\phi^{\ast}\mathcal{O}_{d_{1}}$. Choose sections $\boldsymbol{\mathfrak{s}}^{\pm}_{0}$ of the $\operatorname{\mathcal{L}}^{\pm}$ so that the $(\boldsymbol{\mathfrak{s}}^{\pm}_{0},\boldsymbol{\mathfrak{s}}^{\pm}_{1})$ are pencils for the $(X^{\pm},\operatorname{\mathcal{L}}^{\pm})$ with base loci
$B^{\pm}=\{\boldsymbol{\mathfrak{s}}^{\pm}_{0}=\boldsymbol{\mathfrak{s}}^{\pm}_{1}=0\}\subset Z^{\pm}_{1}.$
By the existence of the projective models, there are holomorphic sections $\boldsymbol{\mathfrak{s}}_{\mathcal{M},0}^{\pm}$ of $\mathcal{O}_{d_{1}}\to\mathbb{CP}^{N}$ such that
$\phi^{\pm}B^{\pm}=\{\boldsymbol{\mathfrak{s}}^{\pm}_{\mathcal{M},1}\oplus\boldsymbol{\mathfrak{s}}^{\pm}_{\mathcal{M},0}=0\}\cap\mathcal{M}.$
Let $\boldsymbol{\mathfrak{s}}^{T}_{\mathcal{M},1}\oplus\boldsymbol{\mathfrak{s}}^{T}_{\mathcal{M},0}$ be a $T\in[-1,1]$ parameter family of holomorphic sections of $\mathcal{O}_{\vec{d}}\oplus\mathcal{O}_{d_{1}}\rightarrow\mathbb{CP}^{N}$ such that $\boldsymbol{\mathfrak{s}}^{\pm 1}_{\mathcal{M},j}=\boldsymbol{\mathfrak{s}}_{\mathcal{M},j}^{\pm}$ for $j=0,1$ and such that the zero loci of the $\boldsymbol{\mathfrak{s}}^{T}_{\mathcal{M},1}$ and $\boldsymbol{\mathfrak{s}}^{T}_{\mathcal{M}}\oplus\boldsymbol{\mathfrak{s}}^{T}_{\mathcal{M},0}$ are transversely cut out for all $T$. The transversality conditions can be satisfied using the fact that failure of transversality is a $\operatorname{codim}_{\mathbb{C}}\geq 1$ condition in the projective spaces of sections $\mathbb{CP}(H^{0}(\mathcal{M},\mathcal{O}_{\vec{d}}\oplus\mathcal{O}_{d_{1}}))$. A parametric Darboux-Moser-Weinstein argument as in the proof of Lemma 2.4.1 gives us symplectomorphism
$\psi:(\phi^{+}Z^{+}_{1},\omega_{\mathcal{O}_{d_{1}}})\rightarrow(\phi^{-}Z^{-}_{1},\omega_{\mathcal{O}_{d_{1}}}),\quad\psi\phi^{+}B^{+}=\phi^{-1}B^{-}.$
So $\phi=(\phi^{-})^{-1}\psi\phi^{+}$ a symplectomorphism between the $(X^{\pm},\omega_{\operatorname{\mathcal{L}}^{\pm}})$ preserving the base loci $B^{\pm}$. ∎
@@ 4.2. Cabling pencil pairs
Theorem 4.2.2 below allows us to determine an infinite family of pencil pairs from a single pencil pair, when the original pencil pair is given by Lemma 4.1.3. The results can be seen as a high-dimensional generalization of Lemma 3.2.2. It’s proof depends on the following general lemma. We suspect that this result is known to experts but were unable to find a proof in the literature.
Theorem 4.2.1.
Let $Z\subset X$ be a divisor of be a very ample line bundle $\operatorname{\mathcal{L}}$ over a closed variety $X$. View $Z$ and $X$ as symplectic manifolds with their symplectic structures inherited from $(\mathbb{CP}^{N},\omega_{\mathbb{CP}^{N}}),N=\dim H^{0}(X,\operatorname{\mathcal{L}})-1$ via the embedding of Equation (2.4.1). For $k\geq 0$, let $Z_{k}$ be a divisor of $\operatorname{\mathcal{L}}^{\otimes k}=\mathcal{O}_{k}|_{X}$ intersecting $Z$ transversely. Then as a symplectic manifold, $Z_{k}\subset X\subset\mathbb{CP}^{N}$ is a $k$-fold branched covering of $Z$ with branch locus $Z\cap Z_{k}$. In particular, the symplectomorphic type of the pairs $(Z_{k},Z\cap Z_{k})$ and $(Z_{k},B_{k})$, where $B_{k}$ is the base locus of a pencil on $X$ containing $Z_{k}$, depends only on $\operatorname{\mathcal{L}}|_{Z}$.
@@
We first explain what we’d like to do and why it wouldn’t work. Assume that $Z$ is cut out by some $\boldsymbol{\mathfrak{s}}_{0}\in H^{0}(X,\operatorname{\mathcal{L}})$. Take $Z_{k}$ to be cut out by a generic section $\mathcal{C}^{1}$-close to $\boldsymbol{\mathfrak{s}}_{0}^{\otimes k}$, so that $Z_{k}$ is contained in a tubular neighborhood $N$ of $Z$ in $X$. Then we can map $Z_{k}$ onto $Z$ by restricting a projection $\pi_{N}:N\to Z$ to $Z_{k}$. If $\pi_{N}$ is holomorphic then $\pi_{N}|_{Z_{k}}$ will be holomorphic, and we can try to prove that $\pi_{N}|_{Z_{k}}$ is a branched covering. Unfortunately, this argument does not work in general as our subvariety $Z\subset X$ is not guaranteed to have a neighborhood $N$ with a holomorphic projection $\pi_{N}$, cf. [MR78]. Instead we’ll deform $Z_{k}\subset X$ to a variety $Z_{k}^{\prime}$ inside the total space $\operatorname{Tot}(\operatorname{\mathcal{L}}|_{Z})$ of the line bundle $\operatorname{\mathcal{L}}|_{Z}\to Z$ by embedding both $X$ and $\operatorname{Tot}(\operatorname{\mathcal{L}}|_{Z})$ into a large projective space. Then $Z^{\prime}_{k}$ and $\operatorname{Tot}(\operatorname{\mathcal{L}}|_{Z})$ have holomorphic projections to $Z$.
Let $\vec{1}=[1:\cdots:1]\in\mathbb{CP}^{N+1}$. Viewing $\mathbb{CP}^{N}=\{z_{N+1}=0\}\subset\mathbb{CP}^{N+1}$, we can think of $Z$ and $X$ as being contained in $\mathbb{CP}^{N+1}$. The map
$\pi_{\vec{1}}:(\mathbb{CP}^{N+1}\setminus\{\vec{1}\})\to\mathbb{CP}^{N},\quad[z_{0}:\cdots:z_{N}:z_{N+1}]\mapsto[z_{0}-z_{N+1}:\cdots:z_{N}-z_{N+1}]$
is a holomorphic submersion using which we can identify $\mathbb{CP}^{N+1}\setminus\{\vec{1}\}$ with the total space of the line bundle $\mathcal{O}_{1}\to\mathbb{CP}^{N}$. The spaces $\pi_{\vec{1}}^{-1}Z$ and $\pi_{\vec{1}}^{-1}X$ are then smooth, quasi-projective varieties in $\mathbb{CP}^{N+1}$. They are projective cones over $Z$ and $X$, with a point $\vec{1}$ removed from each and can be smoothly compactified inside of a blow up at $\mathbb{CP}^{N+1}$ at the point $\vec{1}$.
Without loss of generality, we assume that $Z\subset X$ is the zero locus of $z_{0}$ in $X$. Indeed, a basis $(\boldsymbol{\mathfrak{s}}_{j})$ of $H^{0}(X,\operatorname{\mathcal{L}})$ used to define the embedding of Equation (2.4.1) can be chosen so that $\boldsymbol{\mathfrak{s}}_{0}$ is the section cutting out $Z$. Then we can express $Z$ as $\{z_{N+1}=0\}$ inside of $\pi_{\vec{1}}^{-1}Z$ or as $\{z_{0}=z_{N+1}=0\}$ inside of $\pi_{\vec{1}}^{-1}X$.
For $\boldsymbol{\mathfrak{s}}_{k}$ a $\deg=k$ homogeneous polynomial in variables $z_{1},\cdots,z_{N}$ define
$\sigma_{N}=z_{0}^{k}+\boldsymbol{\mathfrak{s}}_{k}\in H^{0}(\mathbb{CP}^{N},\mathcal{O}_{k}).$
We assume that $\sigma_{N}$ is transversely cut out when viewed as a section of $\mathcal{O}_{k}|_{X}=\operatorname{\mathcal{L}}^{\otimes k}$ and identify its zero locus with the subvariety of interest,
$Z_{k}=\{\sigma_{N}=0\}\cap X.$
As a symplectic manifold we may identify $Z_{k}$ with the zero locus of any transversely cut out section of $\operatorname{\mathcal{L}}^{\otimes k}\to X$ by following the proof of Lemma 2.4.1. We also assume that $Z\cap Z_{k}$ is transversely cut out by $z_{0}\oplus\sigma_{X}\in H^{0}(X,\operatorname{\mathcal{L}}\oplus\operatorname{\mathcal{L}}^{\otimes k})$. This means that $Z\cap Z_{k}$ is cut out by $\boldsymbol{\mathfrak{s}}_{k}$ viewed as an element of $H^{0}(Z,\operatorname{\mathcal{L}})$.
Define a section $\sigma_{N+1}$ of $\mathcal{O}_{k}$ and a subvariety $V_{k}$ of $\pi_{\vec{1}}^{-1}(X)$ as follows:
$\sigma_{N+1}=z_{0}^{k}+\boldsymbol{\mathfrak{s}}_{k}+z_{N+1}^{k}\in H^{0}(\mathbb{CP}^{N+1},\mathcal{O}_{k}),\quad V_{k}=\{\sigma_{N+1}=0\}\cap\pi_{\vec{1}}^{-1}(X).$
Applying variations to $\boldsymbol{\mathfrak{s}}_{k}$, we can assume that $V_{k}$ is transversely cut out of $\pi_{\vec{1}}^{-1}(X)$. Take note of the intersections of $V_{k}$ with the $z_{0}=0$ and $z_{N+1}=0$ hyperplanes:
$\begin{gathered}Z_{k}=V_{k}\cap\{z_{N+1}=0\}\subset X,\quad Z_{k}^{\prime}:=V_{k}\cap\{z_{0}=0\}\subset\pi_{\vec{1}}^{-1}(Z),\\ Z_{k}\cap Z=Z_{k}^{\prime}\cap Z=\{\boldsymbol{\mathfrak{s}}_{k}=0\}\cap Z.\end{gathered}$
We claim that both $Z_{k}$ and $Z_{k}^{\prime}$ are transversely cut out of $V_{k}$. To establish transversality for $Z_{k}$ observe that along $X$, $T\pi_{\vec{1}}^{-1}(X)$ splits as a direct sum of $TX$ and the tangent spaces to the fibers of $\pi_{\vec{1}}^{-1}$. Then $d\sigma_{N+1}|_{X}=d\sigma_{N}$ is non-vanishing along the $TX$ summand and $dz_{N+1}$ is non-vanishing along the complementary summand. Transversality for $Z_{k}^{\prime}$ follows a similar argument by splitting the tangent space $T\pi_{\vec{1}}^{-1}(X)$ over $Z$ as a direct sum of $T\pi_{\vec{1}}^{-1}Z$ and a normal bundle to $Z$ in $X$. It follows that a generic $1$-parameter family of sections of $\operatorname{span}(z_{0},z_{1})\subset H^{0}(\mathbb{CP}^{N+1},\mathcal{O}_{1})$ interpolating between $z_{0}$ and $z_{N+1}$ will cut out a $1$-parameter family of subvarieties of $V_{k}$ which interpolate between $Z_{k}$ and $Z_{k}^{\prime}$. We assume of course that the zero-loci of the hyperplane sections in the $1$-parameter family miss the point $\vec{1}\in\mathbb{CP}^{N+1}$. Through such a $1$-parameter family, every intersection with $Z$ yields $Z_{k}\cap Z$. Again by the logic of Lemma 2.4.1, the pairs $(Z_{k},Z_{k}\cap Z)$ and $(Z_{k}^{\prime},Z_{k}^{\prime}\cap Z)$ are symplectomorphic. To establish that $(Z_{k},B_{k})$ and $(Z_{k}^{\prime},B_{k}^{\prime})$ are symplectomorphic (with $B_{k}$ and $B_{k}^{\prime}$ being the base loci of pencils containing $Z_{k}$ and $Z_{k}^{\prime}$, respectively), we can take another section of $\sigma_{N+1}^{\prime}$ of $\mathcal{O}_{k}$ and keep track of the intersection of the zero locus of $\sigma_{N+1}^{\prime}$ with our subvarieties in the $1$-parameter family.
To complete the proof we must established that $Z^{\prime}_{k}\subset\pi_{\vec{1}}^{-1}(Z)$ is a branched covering of $Z$. This subvariety is cut out of by the section
$\sigma^{\prime}=\boldsymbol{\mathfrak{s}}_{k}+z^{k}_{N+1}\in H^{0}(\pi_{\vec{1}}^{-1}(Z),\mathcal{O}_{k}).$
Identify $\pi_{\vec{1}}^{-1}(Z)$ with the total space $\operatorname{Tot}(\mathcal{O}_{1})=\operatorname{Tot}(\operatorname{\mathcal{L}})$ over $Z$ with projection map $p=\pi_{\vec{1}}$, and $\mathcal{O}_{1}$ as $p^{\ast}(\operatorname{\mathcal{L}}|_{Z})$. With respect to this identification, $z_{N+1}$ is the tautological section $T$ defined by the property that $T(v_{z})=v_{z}$ for $v_{z}\in p^{-1}(z)$. Similarly $\boldsymbol{\mathfrak{s}}_{k}$ is lifted from the base $Z$ of the projection. Therefore we can view $\sigma^{\prime}$ as a section of $(p^{\ast}\mathcal{O}_{1})^{\otimes k}\to\operatorname{Tot}(\mathcal{O}_{1}|_{Z})$ expressible as $T^{\otimes k}+\boldsymbol{\mathfrak{s}}_{k}$.
The remainder of the proof follows a standard holomorphic branched covering construction. See for example the proof of [L04, Proposition 4.1.6]. Consider the map
$\operatorname{pow}_{k}:\operatorname{Tot}(\mathcal{O}_{1}|_{Z})\to\operatorname{Tot}(\mathcal{O}_{k}|_{Z}),\quad\operatorname{pow}_{k}v_{z}=v_{z}^{\otimes k},\quad z\in p^{-1}(z),z\in Z.$
This is a $k$-fold branched covering whose branch locus is the zero section. If $T_{k}$ is the tautological section of $p_{k}^{\ast}\mathcal{O}_{k}$ over $\operatorname{Tot}(\mathcal{O}_{k}|_{Z})$ where $p_{k}:\operatorname{Tot}(\mathcal{O}_{k}|_{Z})\to Z$ is the projection, then $\operatorname{pow}^{k}Z^{\prime}_{k}$ is the zero locus of $\boldsymbol{\mathfrak{s}}_{k}+T_{k}$. This zero locus is exactly the graph of $-\boldsymbol{\mathfrak{s}}_{k}$ in $\operatorname{Tot}(\mathcal{O}_{k}|_{Z})$. Hence the branch locus of $\operatorname{pow}_{k}|_{Z^{\prime}_{k}}$ is exactly $\{\boldsymbol{\mathfrak{s}}_{k}=0\}\cap Z=Z^{\prime}_{k}\cap Z$. Therefore $p_{k}\circ\operatorname{pow}_{k}:Z^{\prime}_{k}\to Z$ is a holomorphic branched covering with branch locus $Z_{k}^{\prime}\cap Z=Z_{k}\cap Z$ as claimed. ∎
Theorem 4.2.2.
@@
Suppose that $(X^{\pm},\operatorname{\mathcal{L}}^{\pm})$ have divisors with a common projective model. Then for all $k\in\mathbb{Z}_{\geq 1}$, $(X^{\pm},(\operatorname{\mathcal{L}}^{\pm})^{\otimes k})$ have divisors with a common projective model, yielding pencil pairs $(X^{\pm},(\operatorname{\mathcal{L}}^{\pm})^{\otimes k},\phi_{k})$ by Lemma 4.1.3.
Proof.
This follows immediately from Theorem 4.2.1 and the definitions involved. The projective models for polarized divisors $(Z^{\pm}_{1},\operatorname{\mathcal{L}}^{\pm})$ are deformation equivalent through complete intersection subvarieties of some $\mathcal{M}\subset\mathbb{CP}^{N}$ for which $\operatorname{\mathcal{L}}^{\pm}$ are pull-backs of the line bundle $\mathcal{O}_{d_{1}}\to\mathbb{CP}^{N}$. Theorem 4.2.1 then provides that for all $k$, pairs of divisors and branch-loci for $(X^{\pm},(\operatorname{\mathcal{L}}^{\pm})^{\otimes k})$ will be symplectomorphic by some $\phi_{k}$. Such $\phi_{k}$ define the desired pencil pairs. ∎